OpenStudy (anonymous):

What is the distance between the point (3, 7) and the midpoint of the segment between the points (3, 7) and (-6, 4)? please show all your work.

OpenStudy (anonymous):

you can find the distance of the two points of the line \[d=\sqrt{(3+6)^{2}+(7-4)^{2}}\] find that and you have \[\sqrt{89} \] divide that by two and you have the distance

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

can you check it again. The answer is 3 sqrt10/2

OpenStudy (anonymous):

did you give me the points correctly? There is no way that is the answer

OpenStudy (anonymous):

yes.I copied & pasted

OpenStudy (anonymous):

What is the distance between the point (3, 7) and the midpoint of the segment between the points (3, 7) and (-6, 4)?

OpenStudy (anonymous):

omg... i added wrong... it was supposed to be \[\sqrt{90}\] \[(\sqrt{9}\sqrt{10})/2\] \[(3\sqrt{10})/2\]

OpenStudy (anonymous):

very sorry.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

please can you explain me how can we divide with 2 after getting the distance . thanks

OpenStudy (anonymous):

well. the midpoint of the line segment in the question is a certain distance. Because you have one of the end point at (3,7) and you have to find the distance from that end point to the midpoint you can just divide the total dstance by 2 to get the length of the new line segment. this only works if you have to find the distance from an end point to the midpoint. It is just a shortcut. If there was another point outside of said line segment you would have to find the midpoint and then use the distance formula in order to find the distance.

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

Sorry.Thanks a lot